Please note: You are viewing the unstyled version of this website. Either your browser does not support CSS (cascading style sheets) or it has been disabled. Skip navigation.

Dispatch 19: Back to "B"

   Print Change text to small (default) Change text to medium Change text to large

Related Multimedia

October 3 slideshow
» View Slideshow

Rick Krishfield

October 3, 2010

It's another frigid morning when we arrive back at Mooring B site to redeploy the mooring that we recovered two days earlier. Besides recovering an ITP and deploying an ITP yesterday, the mooring group had work practically nonstop to refurbish the mooring sphere and instruments for the mooring to be deployed today. Snow and frost covers everything on deck, making for slippery conditions. In addition, the cold further complicates the preparations, because frozen pipes and hoses need to thawed to provide coolant to the mooring winch, and the capstan heads need to be cleared of ice so that the mooring wire doesn't slip on the drums.

Despite the difficulties, the deck crew and WHOI group redeploy the mooring without incident, paying special attention to safety in these severe conditions. Meanwhile, CTD, XCTD, and ice observing studies continue to be carried out around the mooring and buoy operations. Everything has to be carefully orchestrated to ensure that all of the work can be performed in the increasingly shorter days. We've been fortunate to avoid any major storms which would be severely detrimental to completion of the program.

Last updated: October 7, 2019

whoi logo

Copyright ©2007 Woods Hole Oceanographic Institution, All Rights Reserved, Privacy Policy.
Problems or questions about the site, please contact